OpenStudy (anonymous):

If a ball dropped from a tower reaches the ground after 3.5 seconds, what is the height of the tower? Given: g = -9.8 meters/second2

OpenStudy (anonymous):

\[d = -1/2 * a * t^2 = -1/2 * -9.8 * 3.5^2 = 60.025\] So, about 60 meters.

OpenStudy (anonymous):

Thanks a million!

OpenStudy (anonymous):

No problem!